A train crosses a pole at the speed of 40 km per hour in 27 seconds. What is the length of the train ?
Answer by mananth(16946)

1 km/hr
1 km = 1000 m
1 hour = 3600 seconds
1000/3600 m/s
=5/18 m/s
40 km/h = 40 *(5/18) m/s
length of train = distance traveled to cross the pole
d= speed * time
=40 *(5/18)*27
20*5*3
=300 meters